OpenStudy (anonymous):

As Mars revolves around the sun, it travels at a rate of approximately 15 miles per second. Convert this rate to miles per minute. At this rate, how many miles will Mars travel in 5 minutes? Do not round your answers.

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

current rate =\(15\dfrac{miles }{second}\) now just replace second by minute using the following \(1~minute=60~second\\\color{blue}{1second=\dfrac{1}{60}~minute}\)

OpenStudy (anonymous):

so is that the answer?

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

by replacing that u will get rate =\(15\times 60\dfrac{miles}{minute}\) for how many miles in 5 minutes just multiply it by 5 that is \(15\times 60\times 5=?\) miles

OpenStudy (anonymous):

4500

OpenStudy (mathmath333):

yea correct